OpenStudy (anonymous):

Solve 3(2x - 8) = 12 A. x = 6 B. x = 3 C. x = -1 D. x = 0

sam (.sam.):

3(2x-8)=12 2x-8=4 2x=12 x=6

OpenStudy (anonymous):

6x-24=12 6x=36 x=6

OpenStudy (anonymous):

6x-24=12 6x=12+24 6x=36 x=36/6 x=6 (A)